OpenStudy (anonymous):

The posts of a hockey goal are 6 ft apart. A player attempts to score by shooting the puck from a point that is 19 ft from on of the posts and 24 ft from another post. Within what angle, must the shot be made. Express you answer to the nearest degree

zepdrix (zepdrix):

Oh still stuck on this one? :D

OpenStudy (anonymous):

Yea :( I wrote it out last night but it didn't work out

zepdrix (zepdrix):

 Law of Cosines states that:\[\huge c^2=a^2+b^2-2ab \cos C\] Applying this to our problem, we get:\[\huge 6^2=19^2+24^2-2(19)(24)\cos C\] And from here, we need to solve for C.

zepdrix (zepdrix):

Subtracting some stuff... we get. \[\huge 6^2-19^2-24^2=-2(19)(24)\cos C\]And thennnn dividing some stuff, we get.. \[\huge \frac{6^2-19^2-24^2}{-2(19)(24)}=\cos C\]

zepdrix (zepdrix):

Anddddddd simplifying it down gives us, \[\huge \frac{-901}{-912}=\cos C\]Canceling out the negatives gives us,\[\huge \frac{901}{912}=\cos C\]

zepdrix (zepdrix):

Then the last step is a little tricky! :) We can rewrite this expression using the inverse function, or we can simply take the inverse cosine of both sides. Let's try doing that. \[\huge \cos^{-1}\left(\frac{901}{912}\right)=\cos^{-1}(\cos C)\]

zepdrix (zepdrix):

\[\huge \cos^{-1}(\cos C)=C\]

zepdrix (zepdrix):

\[\huge C=\cos{-1}\left(\frac{901}{912}\right)\]Last step is to simply plug this into your calculator, make sure you're in degree mode! :D

OpenStudy (anonymous):

Ohh ok I see where I went wrong here. Well thank you for your help :D